What eqation do I need to use for the sum of page numbers on facing pages is 285, find larger number? Please help
Answer by checkley77(12844)

LET THE PAGE NUMBERS BE X & (X+1)
X+(X+1)=285
2X=285-1
2X=284
X=284/2
X=142 IS THE LEFT HAND PAGE NUMBER
142=1=143 IS THE RIGHT HAND PAGE NUMBER.